Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, August 10, 2006

  • The installation of the new casework, counter tops and plumbing fixtures in the MEBs Gross Anatomy Lab and Swing Lab is approaching completion. We anticipate receiving permission to begin using these spaces sometime tomorrow, 8/11/06, although minor punch list items may still remain.
  • In the MEB Lecture Room A, the contractor is pushing to complete the remaining repair work.
    • The remaining acoustic wall paper and the trim has been installed.
    • The electricians are running the electrical and data connections to the individual table top locations.
    • The podium still needs to be connected to the network.
  • Elevator #1 in the Resource Center is back in operation, following repairs completed today. (This was an elevator previously certified, but that had been out of service for some time.)
    • The contractor is also working on correcting electrical problems in Elevator #4. Once these problems are resolved, the elevator will be certified for regular use.
  • Problems have developed in the testing and operation of the newly installed temporary chiller at the Dental School. The contractor is trouble-shooting the condenser water pumps and the chilled water pumps to determine what additional repairs may be needed. Chiller programming and balancing has been delayed until the condenser pumps & chilled water pump problems are corrected. Additional work at the Dental School includes the following:
    • The contractor is trouble-shooting the loading dock ramp flood gate to determine what needs to be repaired in order to make it operational.
    • The architects and engineers are working on the design for the repairs to the adjacent flood walls.
    • Leak testing continues for the ice water system.
  • The work on the temporary electrical and mechanical repairs at 1542 Tulane is proceeding well.
    • The two chiller pumps have been rebuilt and reinstalled.
    • The new building switch gear is expected to be delivered to the site next Monday.
    • The new wiring was delivered yesterday.
    • The damaged electrical panels have been gutted and cleaned. They are now ready for the new wiring to be installed and connected.
    • The contractor is positioned to begin to restore power to the building.
  • In the Central Utility Plant on the downtown campus, the variable speed drive and filters are on order. The necessary valves are ready to be replaced.
  • A contract to add new lighting to various Student Parking Lots has been awarded. We anticipate that the contract documents will be signed and returned within the next week to 10 days. At that time, the work can begin and demolition will start.